A new cordless phone emits 4.0 mW at 5.8 GHz. The manufacturer claims that the phone has a range of 100 feet.

If we assume that the wave spreads out evenly with no obstructions, what is the electric field strength at the base unit 100 feet from the phone?
Answer #1

distance from the source, r = 100 feet

= 100*0.3048

= 30.48 m

Intensity at the given point, I = Power/Area

= 4*10^-3/(4*pi*30.48^2)

= 3.426*10^-7 W/m^2

Intenisty of light wave, I = Emax^2/(2*c*mue)


==> Emax = sqrt(2*c*I*mue)

= sqrt(2*3*10^8*3.426*10^-7*4*pi*10^-7)

= 0.016 N/c